OpenStudy (anonymous):

Lucy has determined a function f(x) that shows the exponential growth of the number of t-shirts Oliver owns each year. Explain how the f−1(x) can be found and what f−1(300) means.

OpenStudy (anonymous):

is the function f(x) given?

OpenStudy (anonymous):

The -1 means the inverse.

OpenStudy (anonymous):

No, the function is not given. It is more of a generalized question.

OpenStudy (anonymous):

f-1(x) can be found by interchanging x and y and then making y the subject. That is, replace x by y and y by x and then make the new y the subject

OpenStudy (anonymous):

That would find the inverse, or the opposite of the original. So, what would the f-1(300) mean?

OpenStudy (anonymous):

f-1(300) shows how many years Oliver has taken to collect 300 t-shirts

OpenStudy (anonymous):

Thank-you SOO incredibly much. How can I thank-you for helping me through those problems? I there anything I can do besides give you a couple of metals?

OpenStudy (anonymous):

are they equations? as in do they equal something or are they simply expressions?

OpenStudy (anonymous):

if f(x) = 2^x and h(x) = 3^x , then f(x) + h(x) would just be 2^x + 3^x

OpenStudy (anonymous):

you can't simplify that expression if that's what you mean

OpenStudy (anonymous):

You know hat? Nevermind. I can do this one. Just let me know if there is anything I can do for you.

OpenStudy (anonymous):

can I see an example of how you combine an exponential function and a linear one?

OpenStudy (anonymous):

It's just like 4^x + 4x + 4, but nevermind. I got this. I just needed equal variable exponents. Thanks anyway.

OpenStudy (anonymous):

By the way, I thought that you at least deserve to know my name. It's Rose. What's yours?

OpenStudy (anonymous):

I am Ranul :)

OpenStudy (anonymous):

Well, I'm closing the question. You can always message me if you want. I think that I had better log off of here for now, but I get on here at least once a week. So, I'll surely see you later. Thank-you again for all of your help. I hope that you enjoyed this. Have a great day!

OpenStudy (anonymous):

It's real late in the night :/ and it's a school night too. Glad I could help you :) I am off to sleep now. Tc, bye